OpenStudy (anonymous):

Angle ACB is an inscribed angle of circle P. Find x if angle ACB measures x° and arc ACB measures (x + 45)° .

OpenStudy (anonymous):

\[m \angle ACB=\frac{1}{2} [360^0 -(x + 45)°] \rightarrow x =\frac{1}{2} [360^0 -(x + 45)°]\]

OpenStudy (anonymous):

\[2x= [360^0 -x - 45] \rightarrow 3x =315 \rightarrow x = \frac {315}{3} = 105^0\] Thus \[x= 105^0\]

OpenStudy (anonymous):

Thank you! :)